OpenStudy (anonymous):

Solve the quadratic equation. (x - 5)2 = 9

OpenStudy (anonymous):

Expand (x - 5)^2

OpenStudy (anonymous):

these are answer choices A) x = 5 or 3 B) x = 8 or 2 C) x = -5 or 3 Eliminate D) x = -8 or 2

OpenStudy (anonymous):

x^2 - 10x + 25 = 9 x^2 - 10x + 16 = 0 x^2 -2x - 8x + 16 = 0 x (x - 2) -8 (x - 2) = 0 (x - 8) (x - 2) = 0 x - 8 = 0 => x = 8 x - 2 = 0 => x = 2

OpenStudy (anonymous):

So it ix B)